Suppose that the line with equation y=mx+b passes through the points (4,5) and (3,-1) Then m+b=?
Answer by nerdybill(7384)

Suppose that the line with equation y=mx+b passes through the points (4,5) and (3,-1) Then m+b=?
.
m is just the slope between the two points:
m = (y2-y1)/(x2-x1)
m = (-1-5)/(3-4)
m = -6/-1
m = 6
.
to find b, plug one point (4,5) and the slope into equation:
y=mx+b
5=6(4)+b
5=24+b
-19 = b
.
So,
m+b = 6-19 = -13